How many liters of a 5% solution of salt should be added to a 19% solution in order to obtain 616 liters of a 15% solution?
Answer by richwmiller(17219)

a+b=616,
0.05*a+0.19*b=0.15*616
a=616-b
0.05*(616-b)+0.19*b=92.4
30.8-0.05b+0.19*b=92.4
0.14*b=61.6
b=440
a=616-b
a=176 b=440
check
0.05*176+0.19*440=0.15*616
8.8+83.6=92.4
92.4=92.4
ok